### v1

# Exciton Condensation in Landau Levels of Quantum Spin Hall Insulators

Abstract
The interplay of band topology and electron interactions can lead to novel quantum states of matter. In this work, we theoretically study the quantum spin Hall insulator (QSHI) in a perpendicular magnetic field. In the noninteracting case, the QSHI with space inversion and uniaxial spin rotation symmetries undergoes a topological transition into a normal insulator phase at a critical magnetic field $B_{\mathrm{c}}$. The exciton condensation in the lowest Landau levels is triggered by Coulomb interactions in the vicinity of $B_{\mathrm{c}}$ and spontaneously breaks the inversion and the spin rotation symmetries. We propose that the electron spin resonance spectroscopy can directly probe the exciton condensation order. Our results should apply to the InAs/GaSb quantum wells and other QSHIs.
[…]
The excited electron-hole pairs in semiconductors can form bound states called excitons due to the Coulomb interaction. If the binding energy overcomes the band gap, excitons can proliferate and condense at low temperature, forming an exciton insulator (EI) Blatt et al. (1962); Keldysh and Kopaev (1965); Jérome et al. (1967). The EI state is well-established in double-layer semiconductors fabricated with GaAs Eisenstein and MacDonald (2004); Eisenstein (2014) and bilayer graphene Liu et al. (2017); Li et al. (2017) in the quantum Hall regime. The interlayer exciton condensation spontaneously breaks the conservation of interlayer electron number difference, and leads to coherent interlayer tunneling Spielman et al. (2000, 2001), quantized Hall drag resistance Kellogg et al. (2002, 2003); Liu et al. (2017); Li et al. (2017) and dissipationless counterflow currents Kellogg et al. (2004); Tutuc et al. (2004); Li et al. (2017), which are unambiguous evidence of exciton condensation.
The possible EI states in the InAs/GaSb QWs Cheng et al. (1995); Du et al. (2017a); Wu et al. (2019a, b); Wang et al. (2023) and 2D transition-metal dichalcogenides Kogar et al. (2017); Wang et al. (2019); Ma et al. (2021); Jia et al. (2022); Sun et al. (2022) have been intensively studied recently. In double-layer semiconductors with type-II band alignment, the interplay of chiral interlayer hopping triggered by spin-orbit coupling and exciton condensation driven by Coulomb interaction leads to a rich quantum phase diagram, including topological QSHI and QAHI, normal insulator (NI), and various EI states breaking the TRS and/or certain lattice symmetries Naveh and Laikhtman (1996); Pikulin and Hyart (2014); Xue and Macdonald (2018); Blason and Fabrizio (2020); Zhu et al. (2019). However, there has been only indirect evidence of EI states in experiments reporting the electric transport anomalies Jia et al. (2022); Sun et al. (2022); Wu et al. (2019a, b); Wang et al. (2023) and the presumable spectroscopic exciton gap Kogar et al. (2017); Cheng et al. (1995); Du et al. (2017a). More direct probe of the exciton condensation order is still desirable.
In this work, we theoretically study the 2D QSHI in a perpendicular magnetic field. The QSHI is described by the Bernevig-Hughes-Zhang (BHZ) model Bernevig et al. (2006). Even though the TRS is explicitly broken by the magnetic field, the BHZ model has extra space inversion and uniaxial spin rotation symmetries, thus the spin Chern number Sheng et al. (2006) is a well-defined topological invariant, $C_{\mathrm{s}}=(C_{+}-C_{-})/2$, in which $C_{\alpha}$ ($\alpha=\pm$) are the Chern numbers in different spin sectors. The spin Hall conductance at zero temperature is proportional to the spin Chern number and is quantized as $\sigma_{\mathrm{sH}}=C_{\mathrm{s}}\cdot e/(2\pi)$. The QSH state in the BHZ model has $C_{\mathrm{s}}=1$, which is robust in a perpendicular magnetic field up to a critical field strength $B_{\mathrm{c}}$, where the two lowest Landau levels (LLs) intersect König et al. (2007); Chen et al. (2012); Zhang et al. (2014) [see Fig. 1 (a)]. The LL crossing induces a topological quantum phase transition into a NI phase with $C_{\mathrm{s}}=0$ above $B_{\mathrm{c}}$.
Figure 1: (a) The evolution of LLs of the QSHI in a perpendicular magnetic field (black lines). The two lowest LLs (thick black lines) intersect at a critical field $B_{\mathrm{c}}$, marking a topological transition from the QSHI to the NI state. The blue lines depict the quasiparticle energy levels with exciton condensation in the vicinity of $B_{\mathrm{c}}$. (b) The spin Hall conductance $\sigma_{\mathrm{sH}}$ at zero temperature is quantized in the noninteracting case and jumps from $e/(2\pi)$ to zero at $B_{\mathrm{c}}$ (black lines), but it is not quantized in the EI phase (blue line).
The quantum phase diagram is further enriched by the Coulomb interaction. In the vicinity of $B_{\mathrm{c}}$, the energy gap of electron-hole excitations is vanishingly small, thus it can be easily overcome by the exciton binding energy. Therefore, the exciton condensation takes place and an EI phase intervenes between the QSHI and the NI phases at low temperature. The exciton condensation spontaneously breaks the inversion and the spin rotation symmetries, thereby the spin Chern number is not well-defined any more, and the spin Hall conductance $\sigma_{\mathrm{sH}}$ is not quantized in the EI phase. Instead, $\sigma_{\mathrm{sH}}$ continuously decreases from $e/(2\pi)$ to zero [see Fig. 1 (b)]. Moreover, we propose that the exciton condensation order can be directly detected with the electron spin resonance (ESR) spectroscopy.
Topological transition in BHZ model.— The QSHI is described by the four-band BHZ model Bernevig et al. (2006),
$H_{0}=\sum_{\mathbf{k}}\Psi_{\mathbf{k}}^{\dagger}H_{0}(\mathbf{k})\Psi_{\mathbf{k}},$ (1)
in which $\Psi_{\mathbf{k}}=(c_{\mathbf{k}+},d_{\mathbf{k}+},c_{\mathbf{k}-},d_{\mathbf{k}-})^{\mathrm{T}}$ denotes the electron operators in the conduction band ($c_{\mathbf{k}\alpha}$) and the valence band ($d_{\mathbf{k}\alpha}$) with spin index $\alpha=\pm$  [^1] To be precise, neither the spin nor the total angular momentum is conserved in the semiconductor QWs due to the spin-orbit coupling and the lattice anisotropy Bernevig et al. (2006), and the index $\alpha=\pm$ is only a label of the Kramers doublet in each band. Nonetheless, we use the term “spin” for convenience.. In this basis, the Hamiltonian is block-diagonal,
$H_{0}(\mathbf{k})=\begin{pmatrix}h(\mathbf{k})&0\\ 0&h^{*}(-\mathbf{k})\end{pmatrix},$ (2)
and
$h(\mathbf{k})=\begin{pmatrix}\frac{\hbar^{2}\mathbf{k}^{2}}{2m_{\mathrm{e}}^{*}}-\mu_{\mathrm{e}}&w(k_{x}-ik_{y})\\ w(k_{x}+ik_{y})&-\frac{\hbar^{2}\mathbf{k}^{2}}{2m_{\mathrm{h}}^{*}}+\mu_{\mathrm{h}}\end{pmatrix},$ (3)
in which $m_{\mathrm{e}}^{*}$ and $m_{\mathrm{h}}^{*}$ are the effective masses in the conduction band and the valence band, respectively. The chiral form of the interband hopping is dictated by the space inversion and space rotation symmetries Bernevig et al. (2006). The system is in the QSH state for $\mu_{\mathrm{e}}+\mu_{\mathrm{h}}>0$. If the interband hopping is turned off, the system at the charge neutral point is a semimetal with an equal density of electrons and holes, and the density $n_{0}$ is given by $\mu_{\mathrm{e}}+\mu_{\mathrm{h}}=\pi\hbar^{2}n_{0}(1/m_{\mathrm{e}}^{*}+1/m_{\mathrm{h}}^{*})$.
The TRS operator $\mathcal{T}=-i\sigma_{2}\mathcal{K}$ satisfies $\mathcal{T}H_{0}(\mathbf{k})\mathcal{T}^{-1}=H_{0}(-\mathbf{k})$, in which $\sigma_{2}$ is the Pauli matrix acting on the spin indices, and $\mathcal{K}$ is the complex conjugation. Besides the TRS, the BHZ model has a space inversion symmetry and a uniaxial spin rotation symmetry. The space inversion operator $\mathcal{I}=\tau_{3}$ satisfies $\mathcal{I}H_{0}(\mathbf{k})\mathcal{I}^{-1}=H_{0}(-\mathbf{k})$, in which $\tau_{3}$ is the Pauli matrix acting on the band indices, because the two bands are derived from $s$- and $p$-like orbitals, respectively Bernevig et al. (2006). The uniaxial spin rotation is generated by $J_{z}=\sigma_{3}/2$, which satisfies $[J_{z},H_{0}(\mathbf{k})]=0$. In realistic materials, the inversion and the spin rotation symmetries can be explicitly broken by extra asymmetry terms in the Hamiltonian, but these terms are usually much smaller than other relevant energy scales such as the hybridization gap in QSHIs Liu et al. (2008); Liu and Zhang (2013), thus such weak asymmetry can be safely neglected at the leading order. Their effect will be discussed later.
For concreteness, the band parameters of the BHZ model in the following calculations are taken according to experiments in the InAs/GaSb QWs Mu et al. (2016); Han et al. (2019), $m_{\mathrm{e}}^{*}=0.04~{}m_{0}$, $m_{\mathrm{h}}^{*}=0.136~{}m_{0}$, $w=0.8~{}\mathrm{eV}\cdot\text{\AA}$, and $n_{0}=1.63\times 10^{15}~{}\mathrm{m}^{-2}$, in which $m_{0}$ is the free electron mass. These band parameters are highly tunable by changing the QW structure Liu et al. (2008); Liu and Zhang (2013); Du et al. (2017b) and applying gate voltages. The main results in this work are valid as long as the QW is in the inverted regime, and should also apply to other QSHIs.
The perpendicular magnetic field enters the Hamiltonian via the Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}-e\mathbf{A}/\hbar$ and the Zeeman coupling $H_{\mathrm{Z}}=\frac{1}{2}\mu_{\mathrm{B}}B\cdot\mathrm{diag}(g_{\mathrm{e}},g_{\mathrm{h}},-g_{\mathrm{e}},-g_{\mathrm{h}})$, in which $\mu_{\mathrm{B}}$ is the Bohr magneton. The $g$-factors $g_{\mathrm{e}}=11.5$ and $g_{\mathrm{h}}=0.9$ are also taken from the experiments Mu et al. (2016). In the Landau gauge, the vector potential $\mathbf{A}=(0,Bx)$, and the momentum $k_{y}$ is conserved. Moreover, the Hamiltonian has a magnetic translation symmetry generated by $\pi_{x}=k_{x}-eBy/\hbar$, which shifts $k_{y}$ by a constant, $e^{-i\lambda\pi_{x}}k_{y}e^{i\lambda\pi_{x}}=k_{y}-eB\lambda/\hbar$. Let us introduce the ladder operators of LLs in the $k_{y}=k$ subspace,
$a_{k}^{\pm}=\frac{1}{\sqrt{2eB/\hbar}}\big{(}k_{x}\mp i(k-eBx/\hbar)\big{)},$ (4)
and denote the $n$-th LL with spin $\alpha$ in the conduction band and the valence band by $|\mathrm{E}_{\alpha},n\rangle$ and $|\mathrm{H}_{\alpha},n\rangle$, respectively, then the interband hopping term hybridizes $|\mathrm{E}_{+},n\rangle$ with $|\mathrm{H}_{+},n-1\rangle$ ($n\geq 1$) in the spin-up sector, and $|\mathrm{E}_{-},n-1\rangle$ with $|\mathrm{H}_{-},n\rangle$ ($n\geq 1$) in the spin-down sector. The hybridized LL energies are given by
$\epsilon_{n}^{+}=\frac{1}{2}(\xi_{n}^{+}+\zeta_{n-1}^{+})\pm\frac{1}{2}\sqrt{(\xi_{n}^{+}-\zeta_{n-1}^{+})^{2}+8neBw^{2}/\hbar},$ (5)
$\epsilon_{n}^{-}=\frac{1}{2}(\xi_{n-1}^{-}+\zeta_{n}^{-})\pm\frac{1}{2}\sqrt{(\xi_{n-1}^{-}-\zeta_{n}^{-})^{2}+8neBw^{2}/\hbar},$ (6)
in which $\xi_{n}^{\pm}=\frac{\hbar eB}{m_{\mathrm{e}}^{*}}(n+1/2)-\mu_{\mathrm{e}}\mp\frac{1}{2}g_{\mathrm{e}}\mu_{\mathrm{B}}B$ and $\zeta_{n}^{\pm}=-\frac{\hbar eB}{m_{\mathrm{h}}^{*}}(n+1/2)+\mu_{\mathrm{h}}\mp\frac{1}{2}g_{\mathrm{h}}\mu_{\mathrm{B}}B$ are the LL energies in the conduction band and the valence band without hybridization, respectively. The hybridization pushes these LLs away from the Fermi energy at the charge neutral point, which is shown in Fig. 1 (a). However, the two lowest LLs $|\mathrm{E}_{+},0\rangle$ and $|\mathrm{H}_{-},0\rangle$ are not hybridized, $\epsilon_{0}^{+}=\xi_{0}^{+}$ and $\epsilon_{0}^{-}=\zeta_{0}^{-}$. They intersect at a critical magnetic field $B_{\mathrm{c}}$,
$B_{\mathrm{c}}=\frac{\mu_{\mathrm{e}}+\mu_{\mathrm{h}}}{\frac{\hbar e}{2m_{\mathrm{e}}^{*}}-\frac{1}{2}g_{\mathrm{e}}\mu_{\mathrm{B}}+\frac{\hbar e}{2m_{\mathrm{h}}^{*}}-\frac{1}{2}g_{\mathrm{h}}\mu_{\mathrm{B}}}.$ (7)
This level crossing cannot be avoided, because these two lowest LLs carry opposite quantum numbers of the inversion and the spin rotation symmetries, thus they cannot hybridize with each other without breaking these symmetries.
At the charge neutral point, the LL crossing at $B_{\mathrm{c}}$ closes the eletron-hole excitation gap, and induces a topological quantum phase transition from the QSHI to the NI state. Although the TRS is explicitly broken, the inversion and the spin rotation symmetries are preserved, thus the topological transition can be characterized by the spin Chern number. Intuitively, the QSHI at zero magnetic field has $C_{\mathrm{s}}=1$, which persists up to the gap closing at $B_{\mathrm{c}}$ due to its topological robustness. Above $B_{\mathrm{c}}$, all LLs from the valence band are filled while all LLs from the conduction band are empty, thus the system is a NI with $C_{\mathrm{s}}=0$. The spin Chern number can be extracted from the spin Hall conductance at zero temperature calculated with the Kubo formula Qi et al. (2006); Zhang et al. (2014),
$\sigma_{\mathrm{sH}}=-\frac{e\hbar^{2}}{2\Omega}\sum_{mn}\mathrm{Im}\frac{\langle m|\sigma_{3}v_{x}|n\rangle\langle n|v_{y}|m\rangle}{(\epsilon_{m}-\epsilon_{n})^{2}}\big{(}f(\epsilon_{m})-f(\epsilon_{n})\big{)},$ (8)
in which $v_{\alpha}=-i[r_{\alpha},H]$ ($\alpha=x,y$) are the electron velocity operators, and $f(\epsilon_{m})$ is the Fermi distribution function of the eigenstate $|m\rangle$ with energy $\epsilon_{m}$. $\sigma_{\mathrm{sH}}$ is quantized at zero temperature and jumps from $e/(2\pi)$ to zero at $B_{\mathrm{c}}$ as shown in Fig. 1 (b), which unambiguously indicates the topological transition from the QSHI to the NI phase.
Figure 2: (a) Density plot of the exciton condensation order parameter $\Delta$. The dashed line depicts the critical temperature. (b) The hybridized energy levels with the exciton condensation at temperatures from $T=0$ (blue lines) to $T=24~{}\mathrm{K}$ (red lines). The parameters in the Coulomb potential $d=10~{}\mathrm{nm}$ and $\epsilon_{\mathrm{r}}=15$. (c,d) Similar results obtained with a weak inversion asymmetry term $\lambda=0.2~{}\mathrm{meV}$. The EI phase boundary is smeared due to the explicit symmetry breaking. The dashed lines in (d) are the hybridized LLs without exciton condensation.
Exciton condensation in LLs.— The vanishingly small electron-hole excitation gap in the vicinity of $B_{\mathrm{c}}$ is susceptible to exciton condensation. We may consider the Coulomb interaction projected into the two lowest LLs, because all other LLs are pushed far away from the Fermi energy,
$H_{\mathrm{I}}=\sum_{kk^{\prime}qq^{\prime}}h_{k^{\prime}q^{\prime}qk}c_{k^{\prime}}^{\dagger}d_{q^{\prime}}^{\dagger}d_{q}c_{k},$ (9)
in which $c_{k}$ and $d_{k}$ are the electron operators in the two lowest LLs $|\mathrm{E}_{+},0\rangle$ and $|\mathrm{H}_{-},0\rangle$ labeled by the momentum $k_{y}=k$, respectively. The interaction coefficients are given by
$h_{k^{\prime}q^{\prime}qk}=\int\mathrm{d}^{2}\mathbf{r}\mathrm{d}^{2}\mathbf{r}^{\prime}\,U(\mathbf{r}-\mathbf{r}^{\prime})\psi_{k^{\prime}}^{*}(\mathbf{r})\psi_{q^{\prime}}^{*}(\mathbf{r}^{\prime})\psi_{q}(\mathbf{r}^{\prime})\psi_{k}(\mathbf{r}),$ (10)
in which $\psi_{k}(\mathbf{r})$ is the real-space wavefunction in the lowest LL, and $U(\mathbf{r})=\gamma/\sqrt{\mathbf{r}^{2}+d^{2}}$ is the interlayer Coulomb potential, in which $d$ is the interlayer distance, $\gamma=e^{2}/(4\pi\epsilon_{0}\epsilon_{\mathrm{r}})$, and $\epsilon_{0}\epsilon_{\mathrm{r}}$ is the dielectric constant  [^2] See the Supplemental Materials for details.. The coefficients satisfy $h_{k^{\prime}+\kappa,q^{\prime}+\kappa,q+\kappa,k+\kappa}=h_{k^{\prime}q^{\prime}qk}$ due to the magnetic translation symmetry.
The possible exciton condensation in the two lowest LLs is captured by the order parameter $\langle c_{k}^{\dagger}d_{q}\rangle=\Delta\delta_{kq}$, which is assumed to respect both the translation and the magnetic translation symmetries. However, the inversion and the spin rotation symmetries are spontaneously broken by the exciton condensation. The mean-field Hamiltonian is given by $H_{\mathrm{MF}}=H_{\mathrm{LL}}+H_{\mathrm{I,MF}}$,
$H_{\mathrm{LL}}=\sum_{k}(\epsilon_{0}^{+}c_{k}^{\dagger}c_{k}+\epsilon_{0}^{-}d_{k}^{\dagger}d_{k}),$ (11)
$H_{\mathrm{I,MF}}=-\sum_{k}(A^{*}c_{k}^{\dagger}d_{k}+Ad_{k}^{\dagger}c_{k}),$ (12)
in which $A=\sum_{k^{\prime}}h_{kk^{\prime}kk^{\prime}}\Delta$ is also independent of $k$. To solve the mean-field Hamiltonian, we first note that the charge neutrality is maintained by setting the Fermi energy $\mu=(\epsilon_{0}^{+}+\epsilon_{0}^{-})/2$ at any temperature. The self-consistency equation is given by
$\Delta=\frac{A}{2E}\big{(}f(-E)-f(E)\big{)},$ (13)
in which $E=\frac{1}{2}\sqrt{(\epsilon_{0}^{+}-\epsilon_{0}^{-})^{2}+4|A|^{2}}$ is the quasiparticle excitation energy.
The solution of the order parameter $\Delta$ and the energy levels hybridized by the exciton condensation at different temperature and magnetic field are plotted in Fig. 2 (a) and (b), respectively. The exciton condensation takes place in a dome-like region in the vicinity of $B_{\mathrm{c}}$ at low temperature as expected, thus an EI phase intervenes between the QSHI and the NI phases. The LL crossing is avoided due to the symmetry breaking, and the electron-hole excitation gap $2E$ never closes in the EI phase. The spin Chern number is not well-defined in the EI phase due to the symmetry breaking, but the spin Hall conductance $\sigma_{\mathrm{sH}}$ can still be calculated with the Kubo formula (8). As shown in Fig. 1 (b), $\sigma_{\mathrm{sH}}$ at zero temperature is not quantized in the EI phase and continuously decreases from $e/(2\pi)$ to zero.
ESR spectroscopy of exciton condensation.— We propose that the ESR is a direct spectroscopic probe of the exciton condensation order. In ESR experiments, an ac magnetic field is applied to induce resonant transitions of electron states, and the electromagnetic absorption rate is measured. We consider an ac magnetic field also aligned in the perpendicular direction and parallel to the static magnetic field (the Voigt configuration). This configuration respects the space inversion and the uniaxial spin rotation symmetries. While the resonant transition between the two lowest LLs is forbidden by these symmetries, such transition is allowed in the EI phase thanks to the spontaneous symmetry breaking induced by the exciton condensation.
The Zeeman coupling of the two lowest LLs with the ac magnetic field is given by $H_{\mathrm{ac}}=-\frac{1}{2}\mu_{\mathrm{B}}B_{\mathrm{ac}}M_{z}\cos(\omega t)$, in which $B_{\mathrm{ac}}$ is the amplitude of the ac magnetic field, and
$M_{z}=\sum_{k}(g_{\mathrm{e}}c_{k}^{\dagger}c_{k}-g_{\mathrm{h}}d_{k}^{\dagger}d_{k}).$ (14)
The absorption rate Oshikawa and Affleck (2002) $I(\omega)=\frac{1}{8}\omega\mu_{\mathrm{B}}^{2}B_{\mathrm{ac}}^{2}\mathrm{Im}\chi_{zz}(\omega)$. The dynamic magnetic susceptibility $\chi_{zz}(\omega)$ is given by
$\begin{split}\chi_{zz}(\omega)&=-i\int_{0}^{\infty}\mathrm{d}t\,\big{\langle}[M_{z}(t),M_{z}(0)]\big{\rangle}e^{i(\omega+i0^{+})t}\\ &=\sum_{mn}\frac{f(\epsilon_{m})-f(\epsilon_{n})}{\omega+i0^{+}+\epsilon_{m}-\epsilon_{n}}|\langle n|M_{z}|m\rangle|^{2},\end{split}$ (15)
in which $|m\rangle$ denotes the eigenstate of the static Hamiltonian with energy $\epsilon_{m}$.
From the mean-field solution in the EI phase, we find
$I(\omega)\propto\frac{\omega|\Delta|^{2}}{f(-E)-f(E)}\delta(\omega-2E).$ (16)
The resonance frequency is the electron-hole excitation gap $2E$, and the absorption rate is proportional to the exciton condensation order $|\Delta|^{2}$, both of which vary with the temperature and the static magnetic field. Therefore, the emergence of a resonance peak in the vicinity of $B_{\mathrm{c}}$ at low temperature is a direct evidence of the spontaneous symmetry breaking induced by exciton condensation.
Effect of weak inversion asymmetry.— In realistic materials, the inversion and the spin rotation symmetries can be explicitly broken either by the intrinsic bulk and structure asymmetry Liu et al. (2008); Liu and Zhang (2013) or by applying an inplane magnetic field in type-II QWs 22footnotemark: 2, thereby the two lowest LLs are hybridized,
$H_{\mathrm{asy}}=-\sum_{k}(\lambda^{*}c_{k}^{\dagger}d_{k}+\lambda d_{k}^{\dagger}c_{k}).$ (17)
Adding this term to the mean-field Hamiltonian, the self-consistency equation of the exciton condensation is modified as follows,
$\Delta=\frac{A+\lambda}{2E^{\prime}}\big{(}f(-E^{\prime})-f(E^{\prime})\big{)},$ (18)
in which the modified quasiparticle excitation energy $E^{\prime}=\frac{1}{2}\sqrt{(\epsilon_{0}^{+}-\epsilon_{0}^{-})^{2}+4|A+\lambda|^{2}}$. As shown in Fig. 2 (c), the exciton condensation order $\Delta$ is nonzero at any temperature and magnetic field due to the explicit symmetry breaking, thus the EI phase boundary is smeared. Nonetheless, the exciton condensation order is strong only in the vicinity of $B_{\mathrm{c}}$ at low temperature and is vanishingly small elsewhere, thus the feature of the ESR spectroscopy is not significantly affected by the weak inversion asymmetry.
Conclusion.— In summary, we theoretically studied the 2D QSHI with the space inversion and the uniaxial spin rotation symmetries in a perpendicular magnetic field. In the noninteracting case, the system undergoes a topological transition from the QSHI to the NI state at a critical magnetic field $B_{\mathrm{c}}$. The Coulomb interaction triggers the exciton condensation in the vicinity of $B_{\mathrm{c}}$ at low temperature and spontaneously breaks the inversion and the spin rotation symmetries. We propose that the ESR spectroscopy in the Voigt configuration can directly probe the exciton condensation order. Our results should apply to the InAs/GaSb QWs and other QSHIs with (approximate) inversion and spin rotation symmetries.
Acknowledgements.
H.-M.P. thanks Jin-Tao Jin for helpful discussions. L.Z. is grateful to Rui-Rui Du, Zhongdong Han and Tingxin Li for related stimulating collaborations, and to Yuan Li and Jia-Wei Mei for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12174387), the Chinese Academy of Sciences (Nos. XDB28000000, YSBR-057, and JZHKYPT-2021-08), and the Innovative Program for Quantum Science and Technology (No. 2021ZD0302600). L.Z. is also supported by the Xiaomi Foundation.